OpenStudy (anonymous):

Graph the function. Label the vertex and axis of symmetry. y=x^2-5

OpenStudy (campbell_st):

if \[y = ax^2 + bx + c\] then the axis of symmetry is \[x = \frac{-b}{2a}\] in your equation a = 1 and b = 0 substitute to find the axis of symmetry For the vertex, substitute the value of x you found for the axis of symmetry onto the original equation \[y = x^2 - 5\]

OpenStudy (campbell_st):

to find the x- intercepts, or roots you need to let y = 0 and solve for x \[0 = x^2 - 5\]